# Three-dimensional Abelian and nonAbelian gauge Higgs theories

### v1

@@ 3.5 Topological transitions in lattice gauge models
Topological phase transitions are characterized by the divergence of the length scale $\xi$ of the critical modes, as in more conventional cases. However, the absence of a local gauge-invariant order parameter forces one to define the length scale $\xi$ from the behavior of extended objects, for instance Wilson loops in gauge theories. The presence of a divergent length allows one to define a universal critical exponent $\nu$. Other exponents that are typically considered in conventional transitions, such as $\eta$ and $\beta$, are instead not defined in the absence of a local order parameter. Topological transitions are also defined in the quantum setting, where one can use $\nu$ and the dynamic exponent $z$ (defined in terms of the size dependence of the gap at the critical point) to characterize the critical behavior.141414 We mention that topological transitions have been also associated with the breaking of higher-order symmetries, see, e.g., Refs. [236, 237, 238, 239, 240]. While conventional symmetries act on local objects defined on the entire system, higher-order symmetries act on extended objects associated with lower-dimensional sub-manifolds. Topological transitions can be interpreted as due to the spontaneous breaking of such generalized symmetries, generalizing the Landau approach, in which transitions are due to the spontaneous breaking of global symmetries.
@@ 3.5.1 The three-dimensional inverted XY model
The free energy of the IXY model is related by duality to that of the XY model with Villain action [42, 65], which implies that the IXY model undergoes a transition (at $\kappa_{c}=0.076051(2)$ [65, 87]) belonging to the XY universality class, but with inverted high- and low-temperature phases [42].
It is important to stress that the above identification of the universality classes is done using the duality between the gauge and the spin system, and that duality only maps the free energy and the related thermal observables. In particular, the magnetic sector present in the spin universality class has no counterpart in the gauge universality class. Viceversa, the charged sector that characterized the IXY transition, see Sections 5.1.2 and 5.2.2, is not present in the spin model. Thus, the XY FP that controls the behavior of the gauge model differs from the FP that is relevant in the XY model. Indeed, they represent FPs obtained by performing RG transformations on two different classes of Hamiltonians—those with gauge invariance and those with global O(2) symmetry, respectively. Therefore, it is more appropriate to distinguish the IXY or gauge XY universality class from the spin XY universality class. We mention that sometimes the IXY or gauge XY universality class is referred to as XY universality class, see, e.g., Ref. [20]. We will not use here this notation, as we indicate with XY× transitions those transitions that admit an effective O(2) vector LGW description with gauge-dependent order parameter, as described in Sec. 3.3.
@@ 3.5.2 Three-dimensional lattice ${\mathbb{Z}}_{Q}$ gauge theories
The lattice ${\mathbb{Z}}_{Q}$ gauge theories with $Q\geq 2$ are paradigmatic models undergoing finite-temperature topological transitions [20], separating a high-temperature deconfined phase from a low-temperature confined phase, that have been investigated in several works, see, e.g., Refs. [119, 241, 242, 243, 244, 245]. Their lattice Hamiltonian reads [^15] It is interesting to observe that, by means of the Fortuin-Kasteleyn representation [246], the ${\mathbb{Z}}_{Q}$ gauge theory can be extended to any complex value of $Q$, see, e.g., Ref. [247]. Gauge models with real values of $Q$ in the interval [0,1] have been studied, e.g., in Refs.[247, 248].
$H_{{\mathbb{Z}}_{Q}}=-2K\,\sum_{{\bm{x}},\mu>\nu}{\rm Re}\left(\lambda_{{\bm{x}},{\mu}}\,\lambda_{{\bm{x}}+\hat{\mu},{\nu}}\,\bar{\lambda}_{{\bm{x}}+\hat{\nu},{\mu}}\,\bar{\lambda}_{{\bm{x}},{\nu}}\right),\qquad\lambda_{{\bm{x}},\mu}=e^{i2\pi n/Q},\quad n\in 0,1,\ldots,Q-1,$ (20)
@@
Their critical behavior can be inferred by using duality. The ${\mathbb{Z}}_{Q}$ lattice gauge model is dual to a specific $Q$-state clock model [143], with $\mathbb{Z}_{Q}$ spin variables $\exp(i2\pi n_{\bm{x}}/Q)$ (where $n_{\bm{x}}=1,...,Q$) associated with the lattice sites, and a Hamiltonian which is symmetric under global ${\mathbb{Z}}_{Q}$ transformations. The 2-state clock model is equivalent to the Ising model, while the 3-state clock model is equivalent to the 3-state Potts model, which undergoes a first-order transition. The nature of the transitions for $Q\geq 4$ can be inferred by studying the RG flow of the LGW field theory associated with a $\mathbb{Z}_{Q}$-symmetric spin system [249]
@@
where $\varphi({\bm{x}})$ is a complex field. The $Q$-dependent potential has dimension $Q$ and is therefore irrelevant for $Q>4$. In this case we can thus set $v=0$, obtaining the standard U(1)-symmetric $\Phi^{4}$ theory for a complex field. This implies an XY critical behavior and an effective enlargement of the symmetry of the critical modes at the transition: While the theory is ${\mathbb{Z}}_{Q}$-symmetric, critical modes are U(1)-symmetric. For $Q=4$, the $Q$-dependent potential has dimension four and represents a cubic anisotropy. The stable FP of its RG flow is again the XY fixed point with $v=0$ [130, 95, 150, 173, 250, 251].161616 For $Q=4$ the anisotropic interaction with coefficient $v$ appearing in the Hamiltonian (21) gives only rise to scaling corrections. However, due to the small absolute value of the corresponding RG dimension [150, 173, 250] $y_{v}=-0.108(6)$, these scaling corrections decay slowly. Beside the stable FP with $v=0$, the LGW model also admits an unstable FP on the line $w=u-6v=0$, where the LGW Hamiltonian can be written as the sum of two identical LGW Hamiltonians with a real scalar field, so that this FP corresponds to an Ising critical behavior. The parameter $w$ parametrizes a relevant perturbation of the decoupled Ising fixed point, with RG dimension $y_{w}=d-2/\nu_{I}=0.17475(2)$, where [252] $\nu_{I}=0.629971(4)$ is the Ising critical exponent. The standard $\mathbb{Z}_{4}$ clock model can be exactly rewritten as a sum of two Ising models [253, 254]. Thus, it corresponds to the LGW theory with $w=0$ and hence it undergoes an Ising transition. Generic four-state clock models, however, undergo XY transitions.
@@
Using duality, the results for models with global $\mathbb{Z}_{Q}$ symmetry allow us to infer the nature of the transitions in the ${\mathbb{Z}}_{Q}$ gauge models. These arguments predict an Ising transition for $Q=2$, a first-order transition for $Q=3$, and an XY transition for any $Q>4$. For $Q=4$ the ${\mathbb{Z}}_{4}$ gauge model has a transition in the Ising universality class because its partition function can be written as that of two independent ${\mathbb{Z}}_{2}$ gauge models [255, 92] (this is analogous to what happens in the standard ${\mathbb{Z}}_{4}$ clock model, see footnote 16). However, this result is specific of the gauge formulation with Hamiltonian (20). Generic discrete models with ${\mathbb{Z}}_{4}$ gauge symmetry are instead expected to undergo transitions belonging to the XY universality class, or, more precisely, to the IXY universality class.
@@
For $Q\to\infty$ we obtain the compact U(1) gauge theory in which the fields $\lambda_{{\bm{x}},\mu}$ are U(1) phases. This model does not have finite-$K$ transitions because of the presence of topological excitations (monopoles) that keep the system always disordered [256]. The limit $K\to\infty$, $Q\to\infty$ at fixed $K/Q^{2}$ is more interesting [90]. In this case one obtains the IXY model defined in Sec. 3.5.1 with $\kappa=K/Q^{2}$.171717 This result implies that the critical value $K_{c}(Q)$ for the $Q$-state gauge model scales as $K_{c}(Q)\simeq\kappa_{\rm nc}^{(c)}\,Q^{2}$ for large $Q$, where $\kappa_{\rm nc}^{(c)}=0.076051(2)$ is the critical coupling of the IXY model. Using the existing estimates [244, 257] of $K_{c}(Q)$, Ref. [244] obtained $K_{c}(Q)\simeq w\,Q^{2}$ with $w=0.076053(4)$, in excellent agreement with the theoretical prediction, see also Ref. [90]. The IXY model is the relevant one for the behavior of the gauge model for $Q\geq 5$ (and also for $Q=4$ in the generic case): For these values of $Q$ transitions in the ${\mathbb{Z}}_{Q}$ gauge model have the same critical behavior as the transition in the IXY model. It is interesting to give a RG interpretation of these results, considering the large-$Q$ ${\mathbb{Z}}_{Q}$ gauge model as a perturbed IXY model, in which the inverse charge $1/Q$ plays the role of perturbation parameter. The IXY critical behavior of the all gauge models with $Q\geq 4$ indicates that the $1/Q$ perturbation of the IXY critical behavior is irrelevant.
@@
Finally, we mention that the size dependence of the Wilson loops $W_{C}$, defined as the product of the link variables along a closed contour $C$ within a plane, provide a nonlocal order parameter at the topological transition of the 3D lattice ${\mathbb{Z}}_{Q}$ gauge models [119]. Indeed, their asymptotic size dependence for large contours changes at the transition, passing from the small-$K$ area law $W_{C}\sim\exp(-c_{a}A_{C})$ where $A_{C}$ is the area enclosed by the contour $C$ and $c_{a}>0$ is a constant (this behavior can be easily checked by strong-coupling expansion computations for $K\ll 1$ [258, 8]), to the large-$K$ perimeter law $W_{c}\sim\exp(-c_{p}P_{C})$ where $P_{C}$ is the perimeter of the contour $C$ and $c_{p}>0$ is a constant. Analogous behavior can be inferred within 2D quantum formulations of the ${\mathbb{Z}}_{2}$ gauge theory, see, e.g., Ref. [20].
3.5.3 Topological transitions within low-temperature Higgs phases
Topological states and related transitions, such as those arising in 3D ${\mathbb{Z}}_{Q}$ gauge theories, can also be realized in theories with a larger gauge group if the reduced global symmetry of the low-temperature Higgs phase gives rise to an effective breaking of the gauge group to a residual smaller subgroup and, in particular, to the ${\mathbb{Z}}_{Q}$ group. Topological transitions of this kind [20, 96, 259, 98] are realized in various lattice gauge systems, as we shall see in the following. For example, we mention LAH models with compact gauge variables and higher-charge scalar fields, see Sec. 6.1, and SU($N_{c}$) gauge LNAH model with scalar matter in the adjoint representation, see Sec. 9.4.
In this context, the relevant gauge symmetry is the residual gauge symmetry $G_{R}$ of the Higgs phase, which is defined as the group of gauge transformations that leave invariant a representative scalar field in the minimum-potential configuration. The representative scalar field is supposed to have been fixed to a specific value by using both the global and the local gauge symmetry. Note that the group $G_{R}$ does not depend on the chosen representative of the scalar field in the Higgs phase, and it is determined by the global symmetry of the ordered phase. Therefore, it does not represent an additional characterization of the Higgs phase, and, more importantly, it has a gauge-independent relevance. The gauge modes associated with the residual gauge group $G_{R}$ may give rise to gauge-independent topological transitions, as supported by the numerical investigations that we will present in the next sections. More precisely, topological transitions are possible if the gauge theory with gauge group $G_{R}$ (and no scalar fields) has a topological transition, such as the ${\mathbb{Z}}_{Q}$ gauge group in three dimensions.
@@ 3.5.3 Topological transitions within low-temperature Higgs phases
4 Abelian-Higgs gauge field theories
Many collective phenomena in condensed-matter physics [7, 12] are described by effective 3D U(1) gauge models in which scalar fields are coupled with an Abelian gauge field. We mention the transitions in superconductors [18, 19], in quantum SU($N$) antiferromagnets [27, 71, 72, 28, 29, 30, 31, 32, 20], and the unconventional quantum transitions between the Néel and the valence-bond-solid phases in two-dimensional antiferromagnetic SU(2) quantum systems [34, 35, 36, 37, 38, 39, 40, 260], which represent the paradigmatic models for the so-called deconfined quantum criticality [33]. The phase structure and the universal features of the transitions in scalar gauge models have been extensively studied [18, 19, 41, 42, 43, 261, 205, 262, 44, 45, 46, 47, 48, 49, 50, 27, 213, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 33, 67, 68, 69, 263, 70, 71, 72, 73, 74, 34, 75, 35, 36, 264, 190, 265, 76, 77, 266, 78, 267, 191, 37, 28, 29, 30, 268, 192, 38, 39, 40, 79, 31, 269, 80, 260, 32, 81, 82, 83, 84, 194, 20, 208, 195, 85, 86, 270, 271, 87, 88, 89, 90, 91, 200, 92, 93], and particular attention has been paid to the role of the gauge fields and of the related topological features, like monopoles and Berry phases, which cannot be captured by effective LGW theories with gauge-invariant scalar order parameters [2, 95, 14, 33, 20].
In this section we introduce the Abelian-Higgs (AH) field theory (AHFT), or scalar quantum electrodynamics (QED), which describes a $d$-dimensional system of degenerate $N$-component scalar fields minimally coupled with an Abelian U(1) gauge field. By properly choosing the lagrangian potential of the scalar fields, we can define models with different global symmetry: here we discuss SU($N$) and SO($N$) invariant models. We outline the known results for the renormalization-group (RG) flow of the models and the corresponding fixed points, which are expected to describe the critical behavior of transitions in which both scalar and gauge field are critical (GFT transitions in the classification of Sec. 3.1).
4.1 SU($N$)-symmetric Abelian-Higgs field theory

### v3

@@ 3.5 Topological transitions in lattice gauge models
Topological phase transitions are characterized by the divergence of the length scale $\xi$ of the critical modes, as in more conventional cases. However, the absence of a local gauge-invariant order parameter forces one to define the length scale $\xi$ from the behavior of extended objects, for instance Wilson loops in gauge theories. The presence of a divergent length allows one to define a universal critical exponent $\nu$. Other exponents that are typically considered in conventional transitions, such as $\eta$ and $\beta$, are instead not defined in the absence of a local order parameter. Topological transitions are also defined in the quantum setting, where one can use $\nu$ and the dynamic exponent $z$ (defined in terms of the size dependence of the gap at the critical point) to characterize the critical behavior.141414 We mention that topological transitions have also been associated with the breaking of higher-order symmetries, see, e.g., Refs. [267, 268, 269, 270, 271]. While conventional symmetries act on local objects defined on the entire system, higher-order symmetries act on extended objects associated with lower-dimensional submanifolds. Topological transitions can be interpreted as due to the spontaneous breaking of such generalized symmetries, somehow extending the Landau approach in which transitions are related to the spontaneous breaking of global symmetries.
@@ 3.5.1 The three-dimensional inverted XY model
The free energy of the IXY model is related by duality to that of the XY model with Villain action [42, 65], which implies that the IXY model undergoes a transition (at $\kappa_{c}=0.076051(2)$ [65, 87] when setting the temperature $T=1$) belonging to the XY universality class, but with inverted high- and low-temperature phases [42].
It is important to stress that the above identification of the universality classes is done using the duality between the gauge and the spin system, and that duality only maps the free energy and the related thermal observables. In particular, the magnetic sector present in the spin universality class has no counterpart in the gauge universality class. Viceversa, the charged sector that characterizes the IXY transition, see Secs. 5.1.2 and 5.2.2, is not present in the spin model. Thus, the XY FP that controls the behavior of the gauge model differs from the FP that is relevant in the XY model. Indeed, they represent FPs obtained by performing RG transformations on two different classes of Hamiltonians—those with local O(2) and global O(2) symmetry, respectively. Therefore, it is more appropriate to distinguish the IXY or gauge XY universality class from the spin XY universality class. [^15] The IXY or gauge XY universality class is sometimes also referred to as XY universality class in the literature, see, e.g., Ref. [20].
@@ 3.5.2 Three-dimensional lattice ${\mathbb{Z}}_{Q}$ gauge theories
The lattice ${\mathbb{Z}}_{Q}$ gauge theories with $Q\geq 2$ are paradigmatic models undergoing finite-temperature topological transitions [20], separating a high-temperature deconfined phase from a low-temperature confined phase, that have been investigated in several works, see, e.g., Refs. [120, 272, 273, 274, 275, 276, 277, 278]. Their lattice Hamiltonian reads [^16] It is interesting to observe that, by means of the Fortuin-Kasteleyn representation [279], the ${\mathbb{Z}}_{Q}$ gauge theory can be extended to any complex value of $Q$, see, e.g., Ref. [280]. Gauge models with real values of $Q$ in the interval [0,1] have been studied, e.g., in Refs.[280, 281].
$H_{{\mathbb{Z}}_{Q}}=-2K\,\sum_{{\bm{x}},\mu>\nu}{\rm Re}\left(\lambda_{{\bm{x}},{\mu}}\,\lambda_{{\bm{x}}+\hat{\mu},{\nu}}\,\bar{\lambda}_{{\bm{x}}+\hat{\nu},{\mu}}\,\bar{\lambda}_{{\bm{x}},{\nu}}\right),\quad\lambda_{{\bm{x}},\mu}=\exp\left(i\frac{2\pi n_{{\bm{x}},\mu}}{Q}\right),\quad n_{{\bm{x}},\mu}\in 0,1,\ldots,Q-1,$ (20)
@@
The critical behavior can be inferred by using duality. The ${\mathbb{Z}}_{Q}$ lattice gauge model is dual to a specific $Q$-state clock model [144], with $\mathbb{Z}_{Q}$ spin variables $\exp(2\pi in_{\bm{x}}/Q)$ (where $n_{\bm{x}}=1,...,Q$) associated with the lattice sites, and a Hamiltonian which is symmetric under global ${\mathbb{Z}}_{Q}$ transformations. The 2-state clock model is equivalent to the Ising model, while the 3-state clock model is equivalent to the 3-state Potts model, which undergoes a first-order transition. The nature of the transitions for $Q\geq 4$ can be inferred by studying the RG flow of the LGW field theory associated with a $\mathbb{Z}_{Q}$-symmetric spin system [282]
@@
where $\varphi({\bm{x}})$ is a complex field. The $Q$-dependent potential has dimension $Q$ and is therefore irrelevant for $Q>4$. In this case we can thus set $v=0$, obtaining the standard U(1)-symmetric $\Phi^{4}$ theory for a complex field. This implies an XY critical behavior and an effective enlargement of the symmetry of the critical modes at the transition: While the lattice model is ${\mathbb{Z}}_{Q}$-symmetric, critical modes are U(1)-symmetric. For $Q=4$, the $Q$-dependent potential has dimension four and represents a cubic anisotropy. The stable FP of its RG flow is again the XY FP with $v=0$ [131, 96, 179, 204, 157, 170].171717 For $Q=4$ the anisotropic interaction with coefficient $v$ appearing in the Hamiltonian (21) gives only rise to scaling corrections. However, due to the small absolute value of the corresponding RG dimension [179, 204, 157] $y_{v}=-0.108(6)$, these scaling corrections decay slowly. Beside the stable FP with $v=0$, the LGW model also admits an unstable FP on the line $w=u-6v=0$, where the LGW Hamiltonian can be written as the sum of two identical LGW Hamiltonians with a real scalar field, so this FP corresponds to an Ising critical behavior. The parameter $w$ corresponds to a relevant perturbation of the decoupled Ising fixed point, with RG dimension $y_{w}=d-2/\nu_{I}=0.17475(2)$, where [159] $\nu_{I}=0.629971(4)$ is the Ising critical exponent. The standard $\mathbb{Z}_{4}$ clock model can be exactly rewritten as a sum of two Ising models [283, 284]. Thus, it corresponds to the LGW theory with $w=0$ and hence it undergoes an Ising transition. Generic four-state clock models, however, undergo XY transitions.
@@
Using duality, the results for models with global $\mathbb{Z}_{Q}$ symmetry allow us to infer the nature of the transitions in the ${\mathbb{Z}}_{Q}$ gauge models. These arguments predict an Ising transition for $Q=2$, a first-order transition for $Q=3$, and an XY transition for any $Q>4$. For $Q=4$ the ${\mathbb{Z}}_{4}$ gauge model has a transition in the Ising universality class because its partition function can be written as that of two independent ${\mathbb{Z}}_{2}$ gauge models [285, 92] (this is analogous to what happens in the standard ${\mathbb{Z}}_{4}$ clock model, see footnote 17). However, this result is specific of the gauge formulation with Hamiltonian (20). Generic discrete models with ${\mathbb{Z}}_{4}$ gauge symmetry are instead expected to undergo transitions belonging to the gauge XY (IXY) universality class.
@@
For $Q\to\infty$ we obtain the compact U(1) gauge theory in which the fields $\lambda_{{\bm{x}},\mu}$ are U(1) phases. This model does not have finite-$K$ transitions because of the presence of topological excitations (monopoles) that keep the system always disordered [286]. The limit $K\to\infty$, $Q\to\infty$ at fixed $K/Q^{2}$ is more interesting [90]. In this case one obtains the IXY model defined in Sec. 3.5.1 with $\kappa=2K/Q^{2}$.181818 This result implies that the critical value $K_{c}(Q)$ for the $Q$-state gauge model scales as $2K_{c}(Q)\simeq\kappa_{\rm nc}^{(c)}\,Q^{2}$ for large $Q$, where $\kappa_{\rm nc}^{(c)}=0.076051(2)$ is the critical coupling of the IXY model. Using the existing estimates [276, 287] of $K_{c}(Q)$, Ref. [276] obtained $2K_{c}(Q)\simeq w\,Q^{2}$ with $w=0.076053(4)$, in good agreement with the theoretical prediction, see also Ref. [90]. The IXY model is the relevant one for the behavior of the gauge model for $Q\geq 5$ (and also for $Q=4$ in the generic case): For these values of $Q$, transitions in the ${\mathbb{Z}}_{Q}$ gauge model have the same critical behavior as the transition in the IXY model. It is interesting to give a RG interpretation of these results, considering the large-$Q$ ${\mathbb{Z}}_{Q}$ gauge model as a perturbed IXY model, in which the inverse charge $1/Q$ plays the role of perturbation parameter. The IXY critical behavior of all ${\mathbb{Z}}_{Q}$ gauge models with $Q\geq 4$ indicates that the $1/Q$ perturbation of the IXY critical behavior is irrelevant.
@@
Finally, we mention that the Wilson loops $W_{C}$, defined as the product of the link variables along a closed contour $C$ within a plane, provide a nonlocal order parameter for the topological transition of the 3D lattice ${\mathbb{Z}}_{Q}$ gauge models [120]. Indeed, their asymptotic size dependence for large contours changes at the transition, varying from the area law $W_{C}\sim\exp(-c_{a}A_{C})$, where $A_{C}$ is the area enclosed by the contour $C$ and $c_{a}>0$ is a constant, which is valid for small values of $K$, (this behavior can be easily checked in the strong-coupling regime for $K\ll 1$ [288, 8]), to the perimeter law $W_{c}\sim\exp(-c_{p}P_{C})$, where $P_{C}$ is the perimeter of the contour $C$ and $c_{p}>0$ is a constant, which is valid for large values of $K$. An analogous behavior is observed in 2D quantum formulations of the ${\mathbb{Z}}_{2}$ gauge theory, see, e.g., Ref. [20].
3.5.3 Topological transitions in low-temperature Higgs phases
Topological transitions, such as those arising in 3D ${\mathbb{Z}}_{Q}$ gauge theories, also occur in theories with a larger gauge group if the reduced global symmetry of the low-temperature Higgs phase gives rise to an effective breaking of the gauge group to a residual smaller subgroup and, in particular, to the ${\mathbb{Z}}_{Q}$ group. Topological transitions of this kind [20, 97, 289, 99] are realized in various lattice gauge systems, as we shall see in the following. For example, we mention LAH models with compact gauge variables and higher-charge scalar fields, see Sec. 6.1, and SU($N_{c}$) gauge LNAH model with scalar matter in the adjoint representation, see Sec. 9.4.
In this context, the relevant gauge symmetry is the residual gauge symmetry $G_{R}$ of the Higgs phase, which is defined as the group of gauge transformations that leave invariant a representative scalar field in the minimum-potential configuration. The representative scalar field is supposed to have been fixed to a specific value by using both the global and the local gauge symmetry. Note that the group $G_{R}$ does not depend on the chosen representative of the scalar field in the Higgs phase, and it is determined by the global symmetry of the ordered phase. Therefore, it does not represent an additional characterization of the Higgs phase, and, more importantly, it has a gauge-independent relevance. The gauge modes associated with the residual gauge group $G_{R}$ may give rise to gauge-independent topological transitions, as supported by the numerical results that we will present in the next sections. More precisely, topological transitions are possible if the gauge theory with gauge group $G_{R}$ (and no scalar fields) has a topological transition, such as the ${\mathbb{Z}}_{Q}$ gauge model in three dimensions.
@@ 3.5.3 Topological transitions in low-temperature Higgs phases
4 Abelian Higgs gauge field theories
Many collective phenomena in condensed-matter physics [7, 12] are described by effective 3D U(1) gauge models in which scalar fields are coupled with an Abelian gauge field. We mention the transitions in superconductors [18, 19], in quantum SU($N$) antiferromagnets [27, 71, 72, 28, 29, 30, 31, 32, 20], and the unconventional quantum transitions between the Néel and the valence-bond-solid phases in two-dimensional antiferromagnetic SU(2) quantum systems [34, 35, 36, 37, 38, 39, 40, 290], which represent the paradigmatic models for the so-called deconfined quantum criticality [33]. The phase structure and the universal features of the transitions in scalar gauge models have been extensively studied [18, 19, 41, 42, 43, 291, 234, 292, 44, 45, 46, 47, 48, 49, 50, 27, 242, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 33, 67, 68, 69, 293, 70, 71, 72, 73, 74, 34, 75, 35, 36, 294, 220, 295, 76, 77, 296, 78, 297, 221, 37, 28, 29, 30, 298, 222, 38, 39, 40, 79, 31, 299, 80, 290, 32, 81, 82, 83, 84, 224, 20, 237, 225, 85, 86, 300, 301, 302, 87, 88, 89, 90, 91, 229, 92, 93], and particular attention has been paid to the role of the gauge fields and of the related topological features, like monopoles and Berry phases, which cannot be captured by effective LGW theories with gauge-invariant scalar order parameters [2, 96, 14, 33, 20].
In this section we introduce the Abelian Higgs (AH) field theory (AHFT), or scalar QED, which describes a $d$-dimensional system of degenerate $N$-component scalar fields minimally coupled with an Abelian U(1) gauge field. By properly choosing the Lagrangian potential of the scalar fields, we can define models with different global symmetry: here we discuss SU($N$) and SO($N$) invariant models. We outline the known results for the renormalization-group (RG) flow of the models and the corresponding FPs, which are expected to describe the critical behavior of transitions in which both scalar and gauge field are critical (GFT transitions in the classification of Sec. 3.1).
4.1 SU($N$)-symmetric Abelian Higgs field theory